OpenStudy (anonymous):

Find the exact value of sin (arccos (3/5)). Please explain.

zepdrix (zepdrix):

These can be a little tricky. When you see a trig function, the inside represents an angle. When you see an INVERSE trig function, the inside represents the relationship of sides of a triangle. Let's start with the inner part first.\[\large \arccos\left(\frac{3}{5}\right)= \theta\]The 3/5 represents the relationship of sides of on a triangle, we can more easily understand this if we REWRITE the function without the arc,\[\large \cos(\theta)=\frac{3}{5}\]Let's draw a triangle for this,Understand how i labeled the sides?